You bought 300 shares of stock of XOMO at $50 per share. You deposited $10,000 in your account to partly _nance this purchase. The rest was loaned by the broker. What was the margin ratio in the account? A week later, the stock price had risen to $52. What was the margin ratio at this point? Ignore any accumulated interest on the loan.
Answer #1

1.
Old Margin ratio=10000/(300*50)=66.67%

2.
New Margin ratio=(10000+300*(52-50))/(300*52)=67.95%